OpenStudy (anonymous):

the sales force a bussiness consists 10 men and 10 women. a production unit of 7 people are choosen at ramdon. what is the prob it will include 2 men and 5 women?

OpenStudy (anonymous):

there are 10C2 ways of choosing 2 men. there are 10C5 ways of choosing 5 women there are 20C7 ways of choosing 7 people from 20 people. multiply the first two and divide by the second.

OpenStudy (anonymous):

what r u multoplying thou?

OpenStudy (anonymous):

10c2 times 10c5 divided by 20c7

OpenStudy (anonymous):

how do u do that on the calculator a reg one

OpenStudy (anonymous):

105c5?

OpenStudy (anonymous):

ok, 10C2 = 10*9/2! 10C5 = 10*9*8*7*6/5! 20C7 = 20*19*18*17*16*15*14/7! remember that 7! = 7*6*5*4*3*2*1, then do what dhatraditya said

OpenStudy (anonymous):

so u mutlply all those?

OpenStudy (anonymous):

10c2= 45

OpenStudy (anonymous):

10c5= 6048

OpenStudy (anonymous):

10C5 = 252

OpenStudy (anonymous):

55814400 20c7

OpenStudy (anonymous):

so now what?

OpenStudy (anonymous):

6048*252/55814400 is your answer

OpenStudy (anonymous):

1524096

OpenStudy (anonymous):

45*252/77520= 1134/7752 = 567/3876 = 189/1292

OpenStudy (anonymous):

=(10C2*10C5)/(20C7)

OpenStudy (anonymous):

ah, I trusted prettychickas answer than 20c7 is 55814400

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

what r u multuplying everything

OpenStudy (anonymous):

is it a decimal?

OpenStudy (anonymous):

there are 10C2 ways of choosing 2 men. there are 10C5 ways of choosing 5 women there are 20C7 ways of choosing 7 people from 20 people. multiply the first two and divide by the second (10C2*10C5)/(20C7) = 45*252/77520= 1134/7752 = 567/3876 = 189/1292

OpenStudy (anonymous):

so u devide those now?

OpenStudy (anonymous):

It depends, what format is required? fractions are ok, so long as they are reduced.

OpenStudy (anonymous):

ok so

OpenStudy (anonymous):

i dont get it i am getting a decimal